OpenStudy (anonymous):

Which of the following is a point of intersection of the graphs of y= 1/2 sin2x and y= 1/2? (a) (pi/4, 1/2) (b) (pi/3, 1/2) (c) (pi/2, 1/2) (d) (pi, 1/2) (e) (2pi, 1/2)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

set the equations equal to each other and solve for x 1/2 = 1/2 sin 2x 1=sin2x pi/2 = 2x x = pi/4

OpenStudy (anonymous):

thank you